OpenStudy (anonymous):

Find 4x − 2y − z, if x = 2, y = 3, and z = −4

OpenStudy (anonymous):

All you have to do is replace the x y z for 2 3 -4 then solve the equation

OpenStudy (anonymous):

I know but I really gotta get this done

OpenStudy (anonymous):

For example; 2x-5y-z if x =1 y= 4 and z =2 2(1)-5(4)-2 2-20-2 =-20

OpenStudy (anonymous):

4(2)-2(3)-4 8-6-4 = -2

OpenStudy (anonymous):

ugh yeah but im really bad at algebra and have no calculator

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

your answer is -2